OpenStudy (anonymous):

find f '(x) from : f (x) = 3x (x ^ 3-1)

OpenStudy (akshay_budhkar):

12x^3-3 using product rule

OpenStudy (anonymous):

f (x) = 3x (x ^ 3-1) =3x^4 -3x f'(x)=12x-3

OpenStudy (akshay_budhkar):

mark 12x^3 dude

OpenStudy (anonymous):

Or easy to understand is ... multiply it to get f9x0 = 3x^4-3x and then differentiate to get 12x-3. (Remember d/dx(x^n) = nx^(n-1)

OpenStudy (anonymous):

ok... good job mark...